OpenStudy (anonymous):

x-2/8 - 2x+1/12=1/3 solve

OpenStudy (anonymous):

x= -16

OpenStudy (anonymous):

by multiplying all term in 24: 3x-6-4x+2=8 x=-12

OpenStudy (anonymous):

x = -16

OpenStudy (anonymous):

Common Denominator 24

OpenStudy (anonymous):

(3x-6)/24 - (4x+2)/24 = 8/24 3x-6 - 4x - 2 = 8 -x = 16 x = -16

OpenStudy (anonymous):

I think its -12\

OpenStudy (anonymous):

x = -0.5 x-2/8-2x+1/12=1/3 -2/8+1/12=1/3+2x-x -2/8+1/12-1/3=x x = - 0.5 :)